OpenStudy (anonymous):

Please help simplify! -2(x+1)^2 over (x^2+1)^2

OpenStudy (anonymous):

In simplified form it is : -2/(x^2+1)

OpenStudy (zzr0ck3r):

it is already

OpenStudy (anonymous):

=(-2(x+1)^2)/(x^2+1)^2 =-2(x^2+1)/(x^2+1)^2 ****(x^2+1) reduces =-2/(x^2+1)

OpenStudy (anonymous):

Oops, yea zzr0ck3r is right , sorry ...I just noticed wht I was doing wrong .